One hose can fill a small swimming pool in 70 minutes. A larger hose can fill the pool in 55 minutes. How long will it take the two hoses to fill the pool working together?
Answer by nerdybill(7384)

One hose can fill a small swimming pool in 70 minutes. A larger hose can fill the pool in 55 minutes. How long will it take the two hoses to fill the pool working together?
Let x = time (mins) it takes for both
then
x(1/70 + 1/55) = 1
multiplying both sides by (70)(55):
x(55 + 70) = (70)(55)
x(125) = 3850
x = 30.8 minutes
or
x = 30 minutes and 48 seconds